### v1

@@ 3 Necessary and Sufficient Conditions of $t$-Generator QC Codes to be Self-Orthogonal
$,\displaystyle\langle([x^{i_{r}}g_{r}],$$\cdots,[x^{i_{r}}g_{r}]),$
$([x^{j_{s}}g_{s}(\overline{k_{r1}}k_{s(\omega+1)}-\overline{k_{r(\omega+1)}}k_{s1})],\cdots,[x^{j_{s}}g_{s}(\overline{k_{r\omega}}k_{sl}-\overline{k_{rl}}k_{s\omega})])\rangle_{E}=0.$ (3)
By Lemma 2, Eq. (3) holds for all $1\leq r,s\leq t$ if and only if $x^{i_{r}+m-j_{s}}g_{r}\overline{g_{s}}\sum_{u=1}^{\omega}(k_{ru}\overline{k_{s(\omega+u)}}-k_{r(\omega+u)}\overline{k_{su}})\equiv 0\mod{x^{m}-1}$ i.e. $h_{r}\mid\overline{g_{s}}\sum_{u=1}^{\omega}(k_{ru}\overline{k_{s(\omega+u)}}-k_{r(\omega+u)}\overline{k_{su}})$ for any $r,s\in\{1,2,\cdots,t\}$. Due to the skew-symmetry of the symplectic inner product, it is sufficient to restrict our attention to the case of $1\leq r\leq s\leq t$. This completes the proof. ∎
4 A Class of 2-Generator QC Codes
In this section, we present the definition and dimension of a new class of 2-generator QC codes. We also investigate the algebraic structure of their dual codes with respect to the Euclidean, symplectic, and Hermitian inner products, as well as their self-orthogonality and dual-containing properties.
4.1 Definition and Dimension of This Class of Codes
We begin by providing the definition and analyzing their dimension of this class of QC codes.
@@ 4.1 Definition and Dimension of This Class of Codes
(1) It is worth noting that the construction in Definition 4 is more general than those in Galindo2018 Guan2022-1 , and duchao2023 . Specifically, setting $v_{2}(x)=0$ corresponds to the construction in Galindo2018 , $v_{2}(x)=1$ corresponds to that in duchao2023 , and $v_{2}(x)=v_{1}(x)$ corresponds to the construction in Guan2022-1 .
(2) The code $\mathcal{C}$ in Definition 4 is defined over $\mathbb{F}_{q^{2}}$ and is denoted by $\mathcal{C}_{q^{2}}$ when the Hermitian inner product is considered.
@@
Let $q=3$, $m=8$, $g_{1}(x)=x^{3}+x^{2}+x+1$, $g_{2}(x)=x^{6}+2x^{4}+x^{2}+2$, $v_{1}(x)=x^{6}+2x^{4}+2x^{2}+1$, $v_{2}(x)=x$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $C$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([v_{2}(x)g_{2}(x)],[g_{2}(x)])$, $C_{1}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $(0,[g_{2}(x)])$, $C_{2}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([g_{2}(x)],[g_{2}(x)])$, and $C_{3}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([v_{1}(x)g_{2}(x)],[g_{2}(x)])$. Through calculations in Magma, it is found that the parameters of the linear code $C$ are $[16,7,6]_{3}$, the parameters of the linear code $C_{1}$ are $[16,7,4]_{3}$, the parameters of the linear code $C_{2}$ are $[16,7,5]_{3}$, and the parameters of the linear code $C_{3}$ are $[16,7,4]_{3}$.
@@
In Example  2, the code $C$ is derived from Definition 4, while $C_{1}$, $C_{2}$, and $C_{3}$ is given by the construction in Galindo2018 Guan2022-1 , and duchao2023 , respectively. It is evident that the parameters of $C$ are better than those of $C_{1}$, $C_{2}$, and $C_{3}$. This demonstrates that the generalized construction given in Definition 4 is capable of producing more and better codes.
@@
4.2 Necessary and Sufficient Conditions for these Codes to be Self-Orthogonal
In this subsection, we present explicit representations of the necessary and sufficient conditions for the self-orthogonality of these 2-generator QC codes based on the results in Section 3.
@@ 4.2 Necessary and Sufficient Conditions for these Codes to be Self-Orthogonal
The QC code $\mathcal{C}$ in Definition 4 is self-orthogonal with respect to the Euclidean inner product if and only if all of the following conditions are satisfied simultaneously:
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$.
Theorem 6.
The QC code $\mathcal{C}_{q^{2}}$ in Definition 4 is self-orthogonal with respect to the Hermitian inner product if and only if all of the following conditions are satisfied simultaneously:
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$.
Theorem 7.
The QC code $\mathcal{C}$ in Definition 4 is self-orthogonal with respect to the symplectic inner product if and only if all of the following conditions are satisfied simultaneously:
$(1)\enspace h_{1}(x)\mid\overline{g_{1}}(x)(\overline{v_{1}}(x)-v_{1}(x)),$
$(2)\enspace h_{1}(x)\mid\overline{g_{2}}(x)(1-\overline{v_{2}}(x)v_{1}(x)),\enspace$
$(3)\enspace h_{2}(x)\mid\overline{g_{2}}(x)(v_{2}(x)-\overline{v_{2}}(x)),$
where $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$.
4.3 Necessary and Sufficient Conditions for These Codes to be Dual-Containing
In this subsection, by analyzing the structure of the dual code of the QC code in Definition 4 under three types of inner products, we derive necessary and sufficient conditions for dual-containing with respect to each inner product.
@@ 4.3 Necessary and Sufficient Conditions for These Codes to be Dual-Containing
Let $\mathcal{C}$ be the QC code in Definition 4. Then its Euclidean dual code $\mathcal{C}^{\perp_{E}}$ is the QC code generated by $([g_{1}^{\perp}(x)],[-\overline{v}_{2}(x)g_{1}^{\perp}(x)])$ and $([-\overline{v}_{1}(x)g_{2}^{\perp}(x)],[g_{2}^{\perp}(x)])$.
@@
Let $\mathcal{D}$ denote the QC code generated by $([g_{1}^{\perp}(x)],[-\overline{v_{2}}(x)g_{1}^{\perp}(x)])$ and $([-\overline{v_{1}}(x)g_{2}^{\perp}(x)],[g_{2}^{\perp}(x)])$. It is easy to get $\gcd(\overline{v_{1}}(x)\overline{v_{2}}(x)-1,x^{m}-1)=1$ from $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$, which yields that
$\dim(\mathcal{D})=2m-\deg(g_{1}^{\perp}(x))-\deg(g_{2}^{\perp}(x))=\deg(g_{1}(x))+\deg(g_{2}(x)).$
@@
where $a_{1}(x),a_{2}(x)\in\mathcal{R}$. Similarly, let $\mathbf{d}$ be an arbitrary codeword in $\mathcal{D}$ and write it as
$\mathbf{d}=b_{1}(x)([g_{1}^{\perp}(x)],[-\overline{v_{2}}(x)g_{1}^{\perp}(x)])+b_{2}(x)([-\overline{v_{1}}(x)g_{2}^{\perp}(x)],[g_{2}^{\perp}(x)])$
@@
where $b_{1}(x),b_{2}(x)\in\mathcal{R}$. Then the Euclidean inner product between $\mathbf{c}$ and $\mathbf{d}$ is given by
$\langle\mathbf{c},\mathbf{d}\rangle_{E}$ $=\langle[a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[b_{1}(x)g_{1}^{\perp}(x)-b_{2}(x)\overline{v_{1}}(x)g_{2}^{\perp}(x)]\rangle_{E}$
$\quad+\langle[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)],[-b_{1}(x)\overline{v_{2}}(x)g_{1}^{\perp}(x)+b_{2}(x)g_{2}^{\perp}(x)]\rangle_{E}$
@@
$\quad+\langle[a_{1}(x)v_{1}(x)g_{1}(x)],[b_{2}(x)g_{2}^{\perp}(x)]\rangle_{E}+\langle[a_{2}(x)g_{2}(x)],[-b_{1}(x)\overline{v_{2}}(x)g_{1}^{\perp}(x)]\rangle_{E}.$
According to Lemma 3, it can be verified that the above equation is equal to 00. Thus, we have $\mathcal{D}\subseteq\mathcal{C}^{\perp_{E}}$. Furthermore, since $\dim(\mathcal{D})=\dim(\mathcal{C}^{\perp_{E}})$, we conclude that $\mathcal{D}=\mathcal{C}^{\perp_{E}}$, which completes the proof. ∎
@@
The QC code $\mathcal{C}$ in Definition 4 is Euclidean dual-containing if and only if all the following conditions hold simultaneously:
$(1)\enspace g_{1}(x)\mid g_{1}^{\perp}(x)(1+\overline{v_{2}}(x)v_{2}(x)),$
$(2)\enspace g_{1}(x)\mid g_{2}^{\perp}(x)(\overline{v_{1}}(x)+v_{2}(x)),$
$(3)\enspace g_{2}(x)\mid g_{2}^{\perp}(x)(1+\overline{v_{1}}(x)v_{1}(x)).$
@@
[…]

### v2

@@ 3 Necessary and sufficient conditions of $t$-generator QC codes to be self-orthogonal
$\langle([$ $x^{i_{r}}g_{r}],\cdots,[x^{i_{r}}g_{r}]),$
$([x^{j_{s}}g_{s}(\overline{k_{r1}}k_{s(\omega+1)}-\overline{k_{r(\omega+1)}}k_{s1})],\cdots,[x^{j_{s}}g_{s}(\overline{k_{r\omega}}k_{sl}-\overline{k_{rl}}k_{s\omega})])\rangle_{E}=0.$ (3.3)
By Lemma 2, Eq. (3.3) holds for all $1\leq r,s\leq t$ if and only if
$x^{i_{r}+m-j_{s}}g_{r}\overline{g_{s}}\sum_{u=1}^{\omega}(k_{ru}\overline{k_{s(\omega+u)}}-k_{r(\omega+u)}\overline{k_{su}})\equiv 0\mod{x^{m}-1},$
i.e. $h_{r}\mid\overline{g_{s}}\sum_{u=1}^{\omega}(k_{ru}\overline{k_{s(\omega+u)}}-k_{r(\omega+u)}\overline{k_{su}})$ for any $r,s\in\{1,2,\cdots,t\}$. Due to the skew-symmetry of the symplectic inner product, it is sufficient to restrict our attention to the case of $1\leq r\leq s\leq t$. This completes the proof. ∎
Remark 1.
It should be noted that the above result can be viewed as a generalization of the conclusion in [31]. Specifically, Case A in Theorem 1 reduces to Theorem 11 in [31] when $t=1$. Moreover, Ezerman et al., inspired by the idea of the Chinese Remainder Theorem, established necessary and sufficient conditions for quasi-twisted codes (a generalization of quasi-cyclic codes) to be self-orthogonal. However, those conditions are subject to a restriction on the code length $ml$, namely that $\gcd(m,q)=1$ [32].
Example 1.
Let $q=2$ and $m=21$. Choose $g(x)=x^{10}+x^{8}+x^{6}+x^{4}+x^{3}+1$, $v_{1}(x)=x^{3}+x+1$, and $v_{2}(x)=x^{3}+x^{2}$. Then we can construct a 1-generator QC code $\mathcal{C}=\langle([v_{1}(x)g(x)],\ [v_{2}(x)g(x))]\rangle$. Using Magma for computation, we find that $\mathcal{C}$ is a binary linear code with parameters $42,11,16]_{2}$, which is a best-known linear code (BKLC); see [36]. According to Theorem 1, the code $\mathcal{C}$ is Euclidean self-orthogonal.
4 A new class of 2-generator QC codes
In this section, we introduce a new class of 2-generator quasi-cyclic codes with generators $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([v_{2}(x)g_{2}(x)],[g_{2}(x)])$, and determine their dimensions. By applying Theorem 1 from Section 3, we derive explicit necessary and sufficient conditions for these codes to be self-orthogonal with respect to the Euclidean, symplectic, and Hermitian inner products. We further investigate the algebraic structure of their dual codes under these three inner products and examine their dual-containing properties. Finally, we provide a comparative analysis between our results and those reported in related literature.
4.1 Definition and analysis of the new QC codes
@@
It is worth noting that the construction in Definition 4 is more general than those in [24][29], and [34]. Specifically, setting $v_{2}(x)=0$ corresponds to the construction in [24], $v_{2}(x)=1$ corresponds to that in [34], and $v_{2}(x)=v_{1}(x)$ corresponds to the construction in [29]. In addition, the code $\mathcal{C}$ in Definition 4 is defined over $\mathbb{F}_{q^{2}}$ when the Hermitian inner product is considered.
@@
Let $q=3$, $m=8$, $g_{1}(x)=x^{3}+x^{2}+x+1$, $g_{2}(x)=x^{6}+2x^{4}+x^{2}+2$, $v_{1}(x)=x^{6}+2x^{4}+2x^{2}+1$, $v_{2}(x)=x$, we have $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$. Let $C$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([v_{2}(x)g_{2}(x)],[g_{2}(x)])$, $C_{1}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $(0,[g_{2}(x)])$, $C_{2}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([g_{2}(x)],[g_{2}(x)])$, and $C_{3}$ be the QC code generated by $([g_{1}(x)],[v_{1}(x)g_{1}(x)])$ and $([v_{1}(x)g_{2}(x)],[g_{2}(x)])$. Through calculations in Magma, it is found that the parameters of the linear codes $C,C_{1},C_{2}$, and $C_{3}$ are $16,7,6]_{3}$, $16,7,4]_{3}$, $16,7,5]_{3}$, and $16,7,4]_{3}$, respectively.
@@
In Example  2, the code $C$ is derived from Definition 4, while $C_{1}$, $C_{2}$, and $C_{3}$ is given by the construction in [24][29], and [34], respectively. It is evident that the parameters of $C$ are better than those of $C_{1}$, $C_{2}$, and $C_{3}$. This demonstrates that the generalized construction given in Definition 4 is capable of producing more and better codes.
@@
Let $h_{i}(x)=\frac{x^{m}-1}{g_{i}(x)}$, $i=1,2$. Then $\mathcal{C}$ in Definition 4 is self-orthogonal with respect to each of the following inner products if and only if all the corresponding conditions in that case hold simultaneously:
  1. Case A.
    Euclidean inner product.
    $(1.1)\quad h_{1}(x)\mid\overline{g_{1}}(x)\bigl{(}1+v_{1}(x)\overline{v_{1}}(x)\bigr{)},$
    $(1.2)\quad h_{1}(x)\mid\overline{g_{2}}(x)\bigl{(}\overline{v_{2}}(x)+v_{1}(x)\bigr{)},$
    $(1.3)\quad h_{2}(x)\mid\overline{g_{2}}(x)\bigl{(}1+v_{2}(x)\overline{v_{2}}(x)\bigr{)}.$
  2. Case B.
    Hermitian inner product.
    $(2.1)\quad h_{1}^{[q]}(x)\mid\overline{g_{1}}(x)\bigl{(}1+v_{1}^{[q]}(x)\overline{v}_{1}(x)\bigr{)},$
    $(2.2)\quad h_{1}^{[q]}(x)\mid\overline{g_{2}}(x)\bigl{(}\overline{v_{2}}(x)+v_{1}^{[q]}(x)\bigr{)},$
    $(2.3)\quad h_{2}^{[q]}(x)\mid\overline{g_{2}}(x)\bigl{(}v_{2}^{[q]}(x)\overline{v}_{2}(x)+1\bigr{)}.$
  3. Case C.
    Symplectic inner product.
    $(3.1)\quad h_{1}(x)\mid\overline{g_{1}}(x)\bigl{(}\overline{v_{1}}(x)-v_{1}(x)\bigr{)},$
    $(3.2)\quad h_{1}(x)\mid\overline{g_{2}}(x)\bigl{(}1-\overline{v_{2}}(x)v_{1}(x)\bigr{)},$
    $(3.3)\quad h_{2}(x)\mid\overline{g_{2}}(x)\bigl{(}v_{2}(x)-\overline{v_{2}}(x)\bigr{)}.$
Proof.
These results follow directly from Theorem 1. ∎
Lemma 4.
[26] Let $f(x)$, $g(x)$, and $k(x)$ be polynomials over $\mathcal{R}$ (Here the base field is $\mathbb{F}_{q^{2}}$). Then, the following property holds for the Hermitian inner product:
@@
Let $\mathcal{C}$ be the QC code in Definition 4. Then the generators of its dual codes under three types of inner products are given as follows:
  1. Case A.
    The Euclidean dual code $\mathcal{C}^{\perp_{E}}$ is the QC code generated by
    $\bigl{(}[g_{1}^{\perp}(x)],\,[-\overline{v}_{2}(x)g_{1}^{\perp}(x)]\bigr{)}\quad\text{and}\quad\bigl{(}[-\overline{v}_{1}(x)g_{2}^{\perp}(x)],\,[g_{2}^{\perp}(x)]\bigr{)}.$
  2. Case B.
    The Hermitian dual code $\mathcal{C}^{\perp_{H}}$ is the QC code generated by
    $\bigl{(}[g_{1}^{\perp_{H}}(x)],\,[-\overline{v}_{2}^{[q]}(x)g_{1}^{\perp_{H}}(x)]\bigr{)}\quad\text{and}\quad\bigl{(}[-\overline{v}_{1}^{[q]}(x)g_{2}^{\perp_{H}}(x)],\,[g_{2}^{\perp_{H}}(x)]\bigr{)}.$
  3. Case C.
    The symplectic dual code $\mathcal{C}^{\perp_{S}}$ is the QC code generated by
    $\bigl{(}[g_{2}^{\perp}(x)],\,[\overline{v}_{1}(x)g_{2}^{\perp}(x)]\bigr{)}\quad\text{and}\quad\bigl{(}[\overline{v}_{2}(x)g_{1}^{\perp}(x)],\,[g_{1}^{\perp}(x)]\bigr{)}.$
@@
Let $\mathcal{D}_{1}$ denote the QC code generated by $([g_{1}^{\perp}(x)],[-\overline{v_{2}}(x)g_{1}^{\perp}(x)])$ and $([-\overline{v_{1}}(x)g_{2}^{\perp}(x)],[g_{2}^{\perp}(x)])$. It is easy to get $\gcd(\overline{v_{1}}(x)\overline{v_{2}}(x)-1,x^{m}-1)=1$ from $\gcd(v_{1}(x)v_{2}(x)-1,x^{m}-1)=1$, which yields that
$\dim(\mathcal{D}_{1})=2m-\deg(g_{1}^{\perp}(x))-\deg(g_{2}^{\perp}(x))=\deg(g_{1}(x))+\deg(g_{2}(x)).$
@@
where $a_{1}(x),a_{2}(x)\in\mathcal{R}$. Similarly, let $\mathbf{d}_{1}$ be an arbitrary codeword in $\mathcal{D}_{1}$ and write it as
$\mathbf{d}_{1}$ $=b_{1}(x)([g_{1}^{\perp}(x)],[-\overline{v_{2}}(x)g_{1}^{\perp}(x)])+b_{2}(x)([-\overline{v_{1}}(x)g_{2}^{\perp}(x)],[g_{2}^{\perp}(x)])$
@@
where $b_{1}(x),b_{2}(x)\in\mathcal{R}$. Then the Euclidean inner product between $\mathbf{c}$ and $\mathbf{d}_{1}$ is given by
$\langle\mathbf{c},\mathbf{d}_{1}\rangle_{E}$ $=\langle[a_{1}(x)g_{1}(x)+a_{2}(x)v_{2}(x)g_{2}(x)],[b_{1}(x)g_{1}^{\perp}(x)-b_{2}(x)\overline{v_{1}}(x)g_{2}^{\perp}(x)]\rangle_{E}$
$+\langle[a_{1}(x)v_{1}(x)g_{1}(x)+a_{2}(x)g_{2}(x)],[-b_{1}(x)\overline{v_{2}}(x)g_{1}^{\perp}(x)+b_{2}(x)g_{2}^{\perp}(x)]\rangle_{E}$
@@
$+\langle[a_{1}(x)v_{1}(x)g_{1}(x)],[b_{2}(x)g_{2}^{\perp}(x)]\rangle_{E}+\langle[a_{2}(x)g_{2}(x)],[-b_{1}(x)\overline{v_{2}}(x)g_{1}^{\perp}(x)]\rangle_{E}.$ (4.1)
According to Lemma 3, it can be verified that Eq. (4.1) is equal to zero. Thus, we have $\mathcal{D}_{1}\subseteq\mathcal{C}^{\perp_{E}}$. Furthermore, as $\dim(\mathcal{D}_{1})=\dim(\mathcal{C}^{\perp_{E}})$, we conclude that $\mathcal{D}_{1}=\mathcal{C}^{\perp_{E}}$. The other two cases can be proved in a similar manner, and the details are left to the reader. ∎
@@
Let $\mathcal{C}$ be the QC code in Definition 4. Then $\mathcal{C}$ is dual-containing with respect to the following inner products if and only if all the corresponding conditions listed in each case hold simultaneously:
  1. Case A.
    For the Euclidean inner product, the conditions are given by
    $(1.1)\quad g_{1}(x)\mid g_{1}^{\perp}(x)\bigl{(}1+\overline{v_{2}}(x)v_{2}(x)\bigr{)},$
    $(1.2)\quad g_{1}(x)\mid g_{2}^{\perp}(x)\bigl{(}\overline{v_{1}}(x)+v_{2}(x)\bigr{)},$
    $(1.3)\quad g_{2}(x)\mid g_{2}^{\perp}(x)\bigl{(}1+\overline{v_{1}}(x)v_{1}(x)\bigr{)}.$
  2. Case B.
    For the Hermitian inner product, the conditions are given by
    $(2.1)\quad g_{1}(x)\mid g_{1}^{\perp_{H}}(x)\bigl{(}1+v_{2}(x)\overline{v}_{2}^{[q]}(x)\bigr{)},$
    $(2.2)\quad g_{1}(x)\mid g_{2}^{\perp_{H}}(x)\bigl{(}\overline{v}_{1}^{[q]}(x)+v_{2}(x)\bigr{)},$
    $(2.3)\quad g_{2}(x)\mid g_{2}^{\perp_{H}}(x)\bigl{(}1+v_{1}(x)\overline{v_{1}}^{[q]}(x)\bigr{)}.$
  3. Case C.
    For the symplectic inner product, the condition are given by
    $(3.1)\quad g_{2}(x)\mid g_{2}^{\perp}(x)\bigl{(}\overline{v_{1}}(x)-v_{1}(x)\bigr{)},$
    $(3.2)\quad g_{2}(x)\mid g_{1}^{\perp}(x)\bigl{(}1-\overline{v_{1}}(x)v_{2}(x)\bigr{)},$
    $(3.3)\quad g_{1}(x)\mid g_{1}^{\perp}(x)\bigl{(}v_{2}(x)-\overline{v_{2}}(x)\bigr{)}.$
@@
$(1)\enspace g_{1}^{*}(x)\mid\overline{g_{1}^{\perp}}(x)(1+v_{2}(x)\overline{v_{2}}(x)),$
$(2)\enspace g_{1}^{*}(x)\mid\overline{g_{2}^{\perp}}(x)(v_{1}(x)+\overline{v_{2}}(x)),$
$(3)\enspace g_{2}^{*}(x)\mid\overline{g_{2}^{\perp}}(x)(1+v_{1}(x)\overline{v_{1}}(x)).$